OpenStudy (anonymous):

how do the graphs of f(x)=x^2 and g(x)= 3/4x^2 relate?

OpenStudy (amtran_bus):

The easiest way to do this is to graph them. You will find that they are both parabolas, because they have x^2 term.

OpenStudy (igreen):

Graph them here: https://www.desmos.com/calculator @imawesome17

OpenStudy (mathmale):

It'd be even more awesome if you'd please use parentheses around that fraction 3/4, if that's what you originally meant: how do the graphs of f(x)=x^2 and g(x)= 3/4x^2 relate unclear how do the graphs of f(x)=x^2 and g(x)= (3/4)x^2 relate clearer (assuming that your 3/4 is a constant coefficient).